OpenStudy (anonymous):

find the real solutions for the following quadratic equation: x^2+6x+3=0

OpenStudy (anonymous):

\[\large \sin^{-1}(\sin \frac{7 \pi}{5}) \implies \color{green}{ \frac{7 \pi}{5}}\]

OpenStudy (zzr0ck3r):

did they tell you a domain restriction?

OpenStudy (anonymous):

What is the exact value of the following expression: sin^-1 (sin 11pie/8)

OpenStudy (zzr0ck3r):

Looking at the last one, what do you think the answer is?

OpenStudy (zzr0ck3r):

this is a function that is the inside of itself, what happens? f(f^(-1)(x)) = x

OpenStudy (anonymous):

well the anwer is 7pie/5 but it had a range restriction of -pie/2, pie/2

OpenStudy (zzr0ck3r):

does the second one?

OpenStudy (anonymous):

yes it also has the same restrictions

OpenStudy (zzr0ck3r):

the range restriciton makes it "go both ways" so that the inverse is also a fuction

OpenStudy (zzr0ck3r):

then its the same deal

OpenStudy (anonymous):

but according to the exact value it's not right because it can't be bigger than 1

OpenStudy (zzr0ck3r):

o wait I read that wrong 1 sec

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

ok the first one is wrong also, sorry. sin(7pi/5) is in the 3rd quad, but that is not within the domain of sin so when we go back it gives you the angle that is in the 4th quad

OpenStudy (zzr0ck3r):

so sin^(-1)(sin(7pi/5)) = the anlge that is the same y value as 7pi/5 but in the 4th quad

OpenStudy (zzr0ck3r):

make sense?

OpenStudy (zzr0ck3r):

so take 7pi/5 and minus pi from it. this is how much you need to subtract from 0 to get the angle so the answer is -(7pi/5-pi) = sin^(-1)(sin(7pi/5)) = -2pi/5 = aprx -1.25664

OpenStudy (zzr0ck3r):

and 11pi/8 is the same deal sin^(-1)(sin(11pi/8)) = -(11pi/8-pi)

OpenStudy (anonymous):

wow